OpenStudy (anonymous):

It takes Shonda 2/3 of an hour to dry one load of laundry in her dryer. How many hours will it take her to dry 5 loads of laundry?

OpenStudy (sleepyjess):

Welcome to OpenStudy! If it takes 2/3 of an hour to do one load, how long would it take for 2 loads?

OpenStudy (anonymous):

I dont know thats my trouble.

OpenStudy (sleepyjess):

Okay, well let's do it this way, \(\dfrac{2}{3} * \dfrac{2}1\) Now just multiply across the top and bottom.

OpenStudy (sleepyjess):

2*2=? 3*1=?

OpenStudy (anonymous):

4/3

OpenStudy (sleepyjess):

Yes, now how long would it take for 3 loads?

OpenStudy (sleepyjess):

Or just go straight to 5 loads :) \(\dfrac{2}3 * \dfrac 51\)

OpenStudy (anonymous):

So 10/3? Would we simplify it?

OpenStudy (sleepyjess):

Yes, how many times does 10 go into 3?

OpenStudy (anonymous):

none

OpenStudy (sleepyjess):

Whoops, 3 into 10

OpenStudy (anonymous):

ok then 3 haha

OpenStudy (sleepyjess):

Yes, and we'll have 1/3 left over, so 3 1/3 :)

OpenStudy (anonymous):

Thank you!

OpenStudy (sleepyjess):

No problem! Thank you!